Ji-woo conducted a probability experiment. The outcomes were the colours red, blue and green. The probability of red was 2/5 and the probability of green was 1/5. What might her experiment have been? Explain how you decided.
Answer by Edwin McCravy(20054)

Therefore, a likely experiment would be

Ji-woo has a bag of 10 balls, 4 of them red, 2 of them green, and the rest
blue.  She plans to draw a ball from the 10 each day of the month of September,
which has 30 days.  About how many days in September should she theoretically
expect to draw: 

(1) a red ball?  _____
(2) a green ball? _____
(3) a blue ball? _____

She decided this could be the experiment because she would know P(red)=4/10=2/5,
P(green)=2/10=1/5, So P(blue)=1-2/5-1/5=2/5.  And her theoretical expectation
for the number of days in September she would draw each color would be the
30 days multiplied by the probability of each color, or
 
(1) E(red) = (30)(2/5) = 12 red balls
(2) E(green) = (30)(1/5) = 6 green balls
(3) E(blue) = (30)(2/5) = 12 blue balls

Remark:
Since 12 is twice 6, then if she decided to continue the experiment for many
days, she should expect to have gotten about the same total number of red balls
as the total number of blue balls, and about half as many for the total number
of blue balls.  

Edwin